OpenStudy (anonymous):

Maria has three children. There is two years age difference between each child. The total ages of all three children is 36 years. Rosa is the youngest child. How old is rosa? PLEASE NO CONFUSING MATH QUESTIONS JUST PLEASE HELP ME OR DONT BOTHER WRITING I NEED THIS DONE!

OpenStudy (anonymous):

rosa would be 18

OpenStudy (anonymous):

x = youngest child x + 2 = 2nd child x + 4 = 3rd child x + x + 2 + x + 4 = 36 (now combine like terms and solve for x.

OpenStudy (anonymous):

i take my answer back lol kelliegirl33 looks more right then mine

OpenStudy (anonymous):

if I had $1 for every mistake I made, I would be rich...lol

OpenStudy (anonymous):

haha lol me to